# Dynamic Q&A of Clinical Documents with Large Language Models

### v1

@@ Abstract
Electronic health records (EHRs) contain a wealth of valuable patient information embedded within clinical notes. However, the ever-increasing volume and complexity of these unstructured clinical notes makes it extremely challenging for researchers and clinicians to manually locate and extract relevant information in a timely manner. Intelligent systems that can understand clinical documents and enable users to efficiently query and retrieve key details on demand are urgently needed. In this work our aim is to develop a natural language conversational interface powered by state-of-the-art open source large language models (LLMs) that allows users to explore clinical notes through dynamic question-answering. A chatbot interface was built using Langchain, a framework for creating and managing applications powered by large language models, paired with powerful transformer-based language models. The system allows users to simply ask questions in natural language and receive answers extracted from relevant parts of the clinical notes. Extensive experiments were performed to evaluate combinations of semantic embedding models such as the state of the art SentenceTransformers framework, based on the Sentence-BERT embedding models [1], and advanced large language models on their ability to encode both queries and documents for optimal information retrieval. Wizard Vicuna demonstrated the highest accuracy due to its 13B parameters but required heavy compute. To improve inference latency and deployability, model optimization techniques like weight quantization was utilized to reduce latency by approximately 48 times. While results are promising, several limitations remain including model hallucinations, and lack of robust evaluation across diverse medical cases. Addressing these gaps represents important future work to unlock the value in clinical notes and advance clinical decision making through AI.
1 Introduction
@@
Electronic health record (EHR) systems have become ubiquitous in modern healthcare environments [2, 3, 4]. They offer an integrated digital repository of patient health information collected from various clinical encounters across time. EHRs may contain a range of data including demographics, medical and family histories, laboratory test results, radiology images, computed biosignals and crucially, clinical notes generated by physicians, nurses and other providers. These clinical notes, documenting physician-patient interactions, encapsulate a wealth of vital information about the patient’s conditions, treatments, care plans and outcomes in rich natural language. However, a significant limitation of EHRs is that critical patient information is buried within large volumes of unstructured clinical notes that are challenging for providers to efficiently search through and extract relevant details on demand. For instance, a newly assigned physician may need to rapidly comprehend a patient’s care history across multiple specialists to make appropriate decisions. Or a researcher may want to analyze treatment response patterns across populations. These use cases require intelligently sifting through troves of notes to pinpoint key facts, relationships and temporal trends - a formidable manual task. Hence, there is a profound need for intelligent systems that can automatically understand clinical notes written in natural language and enable users to easily query and retrieve targeted information on the fly through conversational interactions.
@@
Traditionally, information retrieval systems have relied on methods like Boolean models, probabilistic models, and term frequency-inverse document frequency (TF-IDF) for surface-level matching of keywords between queries and documents. While these techniques have been useful for basic keyword-based retrieval, they have their limitations. Boolean models utilize operators like AND, OR, and NOT to retrieve documents containing specific keyword combinations, but they lack the ability to understand the nuances and context of natural language queries. Probabilistic models rank results based on keyword occurrences, which can be effective for certain tasks, but they struggle to grasp the underlying meaning of the text. TF-IDF is another widely used method that assigns weights to keywords in queries and documents based on their rarity. This approach helps identify important terms, but it still falls short in capturing the full semantic context and relationships between words. However the primary constraint of these methods is that they lack semantic understanding and the ability to interpret context.
@@
As a result, these traditional techniques often miss relevant information when applied to complex natural language text. They fail to recognize synonyms, antonyms, word variations, and other language intricacies that are essential for accurate and comprehensive information retrieval.
@@
To overcome these limitations and unlock deeper meaning, more advanced techniques such as natural language methods based on large language models (LLMs) [5] can be utilized. The rapidly advancing LLMs, like GPT-3 and its successors, have revolutionized the field of natural language processing, with their contextual understanding, generative capabilities, and fine-tuning adaptability. Integrating LLMs into information retrieval systems holds the potential to deliver more accurate, insightful, and context-aware results, transforming the way users access information in a more personalized and efficient manner.
@@
In this project we explore the utility of modern natural language processing (NLP) methods based on large language models (LLMs) to develop a conversational agent or chatbot that allows users to explore clinical notes through dynamic question answering. By interfacing with the chatbot in natural language as one would with a human assistant, users can extract relevant details buried within notes, facilitating improved clinical care and research.
Contributions: To this end, our key contributions are as follows:
@@
  • We developed a conversational question answering system based large language models that allows natural language interaction for exploring clinical notes
  • We discussed detailed step by step pipeline of the question answering system based on a medical dataset
@@
  • We conducted extensive experiments evaluating combinations of different embedding models and LLMs to determine optimal configurations for accuracy and speed
  • We showed how the latency and deployability of these models can be optimized by using weights quantization
@@
  • We highlighted the superiority of retrieval augmented generation over domain-specific fine-tuning, demonstrating its advantages in our study
  • We highlighted key limitations around model performance, evaluation rigor, and real-world deployment that open the doors for future research in that domain
@@
Organization: The paper is organized as follows: We discuss related works in section 2. Section 3 discusses the approach. Details of the dataset are discussed in section 4. Evaluation and results are discussed in section 5. We discuss the limitations and future work in section 6 and section 7 concludes the paper.
2 Related Work
2.1 LLMs in information retrieval
@@
The intersection of Large Language Models (LLMs) and Information Retrieval (IR) has emerged as a pivotal area of research, transforming traditional approaches and reshaping the landscape of information access.
The work by Zhu et al. [6] provides a strong foundation by offering a comprehensive survey of the diverse applications of LLMs in different facets of IR, including query rewriting, retrieval, re-ranking, and reading. While their survey spans a wide array of studies, our research takes a specific angle, honing in on the integration of LLMs in the information retrieval process. Expanding upon this foundation, we delve into the potential of enhancing traditional retrieval methods by incorporating text generation capabilities. Our aim is to provide a more natural and conversational interaction between users and the system, ensuring a seamless and user-friendly experience.
@@
Another influential contribution in the realm of information retrieval is the tutorial presented by Begoli et al. [7]. This tutorial focuses on advancing both the theoretical and practical aspects of transformer-based approaches, particularly with an emphasis on efficiency and robustness in large-scale information retrieval scenarios. The tutorial aims to equip the audience with insights gained from developing and fine-tuning transformer/BERT-based models for IR tasks. In contrast, our approach focuses on the utilization of off-the-shelf models, enabling the direct retrieval of information from provided resources without the need for intricate fine-tuning procedures. This streamlined method significantly boosts usability and efficiency, circumventing the time and computational expenses associated with fine-tuning. Furthermore, our research demonstrates that the accuracy achieved through our direct utilization method surpasses that attained through domain-specific fine-tuning processes.
2.2 Medical question answering using LLMs
In recent years, the field of medical question answering has witnessed remarkable advancements, driven by the integration of Large Language Models (LLMs) into the domain. The quest for expert-level performance in this grand challenge has seen notable contributions from various research works.
Lee et al. [8] address the unique challenges posed by biomedical text mining. Recognizing the limitations of applying general NLP advancements to biomedical corpora, they introduce BioBERT, a domain-specific language representation model pre-trained on large-scale biomedical corpora. BioBERT outperforms generic models like BERT and previous state-of-the-art models in various biomedical text mining tasks, including biomedical named entity recognition, biomedical relation extraction, and biomedical question answering.
Another work by Chakraborty et al. [9]emphasizes the urgency of accelerating biomedical research in light of the SARS-CoV-2 pandemic. Introducing BioMedBERT, a pre-trained model on the extensive BREATHE v1.0 dataset, the authors showcase its efficacy in Question-Answering (QA) and Information Retrieval (IR) tasks. The model achieves state-of-the-art results on BioASQ datasets and exhibits superior performance when coupled with Elasticsearch for IR tasks on the BioASQ dataset.
Singhal et al. [10] present Med-PaLM 2, a groundbreaking development in medical question answering. Addressing the inherent challenges in prior models, Med-PaLM 2 leverages a combination of improvements in base LLMs (specifically PaLM 2), medical domain fine-tuning, and innovative prompting strategies, including a novel ensemble refinement approach. The model achieved an impressive score of up to 86.5% on the MedQA dataset, surpassing its predecessor by over 19% and establishing a new state-of-the-art.
@@ 2.2 Medical question answering using LLMs
In contrast to the aforementioned works, our work diverges by placing a specific emphasis on the integration of retrieval augmented generation for medical question answering. While the discussed models, such as Med-PaLM 2, BioBERT, and BioMedBERT, have made significant strides in achieving expert-level performance through advancements in base LLMs, domain fine-tuning, and innovative prompting strategies, my approach focuses on enhancing deployability and mitigating computational and time constraints. The integration of retrieval augmented generation aims to strike a balance between performance and resource efficiency, making the model more accessible for widespread adoption in medical settings.
2.3 Impact of embedding models on output of large language models
@@
The exploration of word embeddings and their influence on the output of large language models (LLMs) is an under-explored but crucial aspect in natural language processing (NLP). While numerous studies focus on the applications and performance of LLMs, the specific impact of embedding models on their output remains relatively uncharted territory.
@@
The only relevant work in this area is the work by Asudani et al. [11]. In this work the authors focuses on the application of word embeddings and deep learning models in the context of text analytics tasks. Through a systematic literature review, the authors classify and analyze relevant articles, highlighting the increasing interest in employing these techniques for analysis and prediction. The study categorizes literature based on text analytics applications, encompassing areas like text classification, sentiment analysis, named entity recognition (NER), recommendation systems, biomedical text mining, and topic modeling. The paper concludes that utilizing domain-specific word embeddings and the LSTM model can significantly enhance overall text analytics task performance. In our investigation, we direct our attention to a distinct domain that remains unexplored in the aforementioned study — medical question answering. Furthermore, we delve into a different suite of embedding models and language models, areas that Asudani et al. did not specifically address in their research.
3 Approach
Figure 1: Information retrieval pipeline
@@ 3 Approach
Our approach primarily utilizes Retrieval-Augmented Generation (RAG) [12], a technique that merges the power of pre-trained language models with the benefits of external knowledge sources. RAG employs a two-step process: a retrieval step that fetches relevant documents from a vast corpus of text, and a generation step that crafts responses based on these retrieved documents and the given input query.
@@
This technique is particularly advantageous when the required response includes information that the language model has not encountered during its training. By combining the ability to generalize from language model pretraining with the capacity to retrieve specific details from external sources, RAG can generate highly detailed and accurate responses while avoiding the overhead of fine tuning the language model on specific datasets.
@@
    • Content Splitting: The loaded content is then split into chucks. This step is crucial as it breaks down the text into smaller units, to comply with the embedding models context size limitations. Throughout this step, we intentionally allow for some overlap between the different chunks. This strategy helps to maintain the contextual integrity of the data.
    • Text Embedding: Each chunk of text is then transformed into an embedding. These embeddings are numerical representations that encapsulate the semantic meaning of the text. The transformation process is carried out by a distinct embedding model, the details of which will be discussed later.
    • Vector Store: The embeddings, along with some reference to the original content the embedding was created from, are subsequently stored in a vector database or vector store, which forms the backbone of our retrieval system.
  2. 2.
@@
    • Output Generation: The re-ranked vectors, which represent the selected documents, along with the input prompt, are passed to a language model. This model is responsible for text summarization and final output generation, and it is distinct from the earlier mentioned embedding model.
@@
Table 1: Embedding vs Language models
To achieve the previously described pipeline we utilized a tool called Langchain. LangChain is a powerful and versatile Python framework designed to facilitate the development of state-of-the-art language models and their applications. It offers a wide range of pre-configured models and utilities that allow researchers and developers to focus on the creative aspects of their projects, rather than spending time on repetitive and time-consuming tasks such as model configuration, training, and deployment. With its modular and flexible design, LangChain allows users to easily integrate various components and adapt them to their specific needs.
@@
In our approach, we leveraged multiple components of the LangChain framework. First, we used the Document Store component for efficiently storing and retrieving documents. This component allowed us to create a robust retrieval system, which is a crucial aspect of our RAG model. The Document Store uses a vector database to store embeddings and references to the original content, providing a fast and scalable solution for storing and retrieving large volumes of data.
Second, we employed LangChain’s Retriever component for document retrieval. The retriever uses an index, a data structure designed for quick retrieval, to scan the vector store. Using cosine similarity, it identifies and returns the most relevant documents to the user’s input query. This component’s ability to quickly and accurately retrieve relevant documents is essential for the successful application of our RAG model.
We also integrated LangChain’s Memory component into our system. In the context of language model applications, especially those involving conversational interfaces, the ability to refer to information from earlier interactions is crucial. This is where the Memory component comes into play. The Memory component in LangChain offers utilities for storing information about past interactions, acting like a conversation history that can be read from and written to. By integrating this component into our system, we have not only improved the quality of the responses generated but also ensured a more natural and human-like conversational experience.
@@
Lastly, we implemented the Chain component provided by LangChain, which was instrumental in developing our more complex system requirements. Chains allowed us to create coherent workflows, linking multiple components together. For instance, we created a chain that takes user input, formats it with a specific template, and then passes the formatted response to the language model for generation.
4 Dataset
The Medical Information Mart for Intensive Care (MIMIC-IV) [13] , is a single-center dataset encompassing information about over 40k patients admitted to intensive care units (ICUs) at at the Beth Israel Deaconess Medical Center (BIDMC). This dataset includes a comprehensive range of data such as medications, laboratory measurements, observations, vital signs, care provider notes, procedure codes, fluid balance, diagnostic codes, hospital length of stay, imaging reports, survival data, and more. MIMIC-IV serves a vital role in various applications, supporting industrial research, academic endeavors, and initiatives for quality improvement.
In addition to the MIMIC-IV dataset,we also utilized the MIMIC-IV-Note [14], which offers deidentified free-text clinical notes for patients included in the MIMIC-IV clinical database. MIMIC-IV-Note contains over 331k discharge summaries. It also includes over 2 million radiology reports. These notes have undergone deidentification, ensuring the removal of protected health information (PHI) in accordance with the Health Insurance Portability and Accountability Act (HIPAA) Safe Harbor provision. MIMIC-IV-Note allows researchers to study clinical natural language processing and related areas by providing context to the clinical data within MIMIC-IV.
Just like their predecessors, MIMIC-IV and MIMIC-IV-Note are freely available, but accessing them requires undergoing CITI training. This training provides materials on maintaining patient privacy, handling medical records responsibly, and importantly, it includes the necessary Health Insurance Portability and Accountability Act (HIPAA) training, as the data originates from the United States.
The dataset itself is substantial and comes in CSV files, along with SQL scripts that facilitate loading the data into the RDS (Relational Database Service). This allows for easy exploration, querying, and filtering of the data to prepare it for the large language models. High-level statistics on the dataset are available in Table 2.
Figure 2 depicts the number of notes taken per year, with the years projected into the future as part of the deidentification process. Additionally, Figure 3 displays statistics of the note lengths, ranging from a minimum of 3 characters to a maximum of 60,381 characters.

### v2

# Dynamic Question-Answering of Clinical Documents using Retrieval Augmented Generation

@@ Abstract.
Electronic health records (EHRs) contain a wealth of valuable patient information embedded within clinical notes. The ever-increasing volume and complexity of these unstructured clinical notes however, make it extremely challenging for researchers and clinicians to manually locate and extract relevant information in a timely manner. Intelligent systems that can understand clinical documents and enable users to efficiently query and retrieve key details on demand are urgently needed. In this work we present a natural language conversational interface powered by state-of-the-art open source large language models (LLMs) that allows users to explore clinical notes through dynamic question-answering. We built a chatbot interface using Langchain, a framework for creating and managing applications powered by large language models, paired with powerful transformer-based language models. The system allows users to simply ask questions in natural language and receive answers extracted from relevant parts of the clinical notes. We evaluated combinations of semantic embedding models (EM) such as the state of the art SentenceTransformers framework, based on the Sentence-BERT EM(Reimers and Gurevych, 2019), and advanced large language models on their ability to encode both queries and documents for optimal information retrieval. Wizard Vicuna demonstrated the highest accuracy due to its 13B parameters but required heavy compute capabilities. To improve inference latency and deployability, we used model optimization techniques like weight quantization to reduce latency by approximately 48 times. While results are promising, several limitations remain including model hallucinations, and lack of robust evaluation across diverse medical cases. Addressing these gaps represents important future work to unlock the full potential of clinical notes and advance clinical decision making through artificial intelligence (AI).
Information Retrieval, Question and Answering, Natual Language Processing, Large Language Mo
$c$s: Information systems Question answering
1. Introduction
@@
Electronic health record (EHR) systems have become ubiquitous in modern healthcare environments (King et al., 2014; Manuel et al., 2010a, b). They offer an integrated digital repository of patient health information collected from various clinical encounters across time. EHRs may contain demographics, medical and family histories, laboratory test results, radiology images, computed biosignals and crucially, clinical notes generated by physicians, nurses and other care team members. These clinical notes, documenting interactions with the healthcare system, encapsulate a wealth of vital information about patients’ conditions, treatments, care plans and outcomes in rich natural language. However, a significant limitation of EHRs is that critical patient information is buried within large volumes of unstructured clinical notes that are challenging for providers to efficiently search through and extract relevant details on demand (Sedlakova et al., 2023, 2022). For instance, a newly assigned physician may need to rapidly comprehend a patient’s care history across multiple specialists to make appropriate decisions (Hampton et al., 1975), or a researcher may want to analyze treatment response patterns across populations (doi, 2020). These use cases require intelligently sifting through troves of notes to pinpoint key facts, relationships and temporal trends - a formidable manual task. Hence, there is a profound need for intelligent systems that can automatically understand clinical notes written in natural language and enable users to easily query and retrieve targeted information on the fly through conversational interactions.
@@
Traditionally, information retrieval systems have relied on methods like regular expression matching (Ilie, 2016), Boolean logic (Melucci, 2009), probabilistic models (Amati and Van Rijsbergen, 2002), and term frequency-inverse document frequency (TF-IDF) (Sammut and Webb, 2010) for surface-level matching of keywords between queries and documents. While these techniques have been useful for basic keyword-based retrieval, they have their limitations. Boolean models (Melucci, 2009) utilize operators like AND, OR, and NOT to retrieve documents containing specific keyword combinations, but they lack the ability to understand the nuances and context of natural language queries. Probabilistic models (Amati and Van Rijsbergen, 2002) rank results based on keyword occurrences, which can be effective for certain tasks, such as information retrieval in search engines (Sontag et al., 2012), spam email detection (Lau et al., 2012), and sentiment analysis in text data (Celikyilmaz et al., 2010), but they struggle to grasp the underlying meaning of the text. TF-IDF (Sammut and Webb, 2010) is another widely used method that assigns weights to keywords in queries and documents based on their rarity. This approach helps identify important terms, but it still falls short in capturing the full semantic context and relationships between words. The shared limitations across these approaches are the lack of semantic understanding and the ability to interpret the context holistically.
@@
As a result, these traditional techniques often miss relevant information when applied to complex natural language text. They fail to recognize synonyms, antonyms, word variations, context, and other language intricacies that are essential for accurate and comprehensive information retrieval.
@@
Large language models (LLMs) based on the transformer architecture (Vaswani et al., 2017) can be used to overcome these limitations and unlock the deeper meaning. The rapidly advancing LLMs, like GPT-4 (OpenAI, 2023), have revolutionized the field of natural language processing, with their contextual understanding, generative capabilities, and fine-tuning adaptability (Zhao et al., 2023; Hadi et al., 2023). Integrating LLMs into information retrieval systems holds the potential to deliver more accurate, insightful, and context-aware results, transforming the way users access information in a more personalized and efficient manner.
Embedding models (EMs) serve as powerful tools in natural language processing by representing words or phrases in a continuous vector space where the geometric distances between vectors capture semantic relationships between words. EMs, such as word embeddings (Almeida and Xexéo, 2023), play a crucial role in capturing intricate relationships and contextual nuances within the text.
Retrieval augmented generation (RAG) refers to a novel technique where a model first retrieves relevant information from a large corpus or dataset and then generates responses or outputs based on the retrieved information. This approach combines the benefits of both retrieval-based and generation-based methods, allowing for more contextually relevant and informative responses in conversational systems. Recent research works are exploring the trade-offs between this method and traditional fine-tuning approaches for language models, with some favoring RAG for certain contexts (Greyling, 2021; Balaguer et al., 2024).
@@
In this work, we explore the utility of modern natural language processing (NLP) methods based on a combination of Large Language Models (LLMs), Retrieval Augmented Generation (RAG), and embedding models to develop a conversational agent that enables users to interactively explore clinical notes through dynamic question answering. By ’dynamic question answering,’ we mean that users can engage with the system in a flexible manner, posing inquiries about clinical notes in natural language. By interfacing with the agent akin to interacting with a human assistant, users can extract pertinent details buried within notes, thereby facilitating improved clinical care and research endeavors.
Our work serves as a demonstration of practical applicability, particularly in information extraction from clinical notes and the development of chatbots utilizing this documentation. Highlighting its relevance in assisting users or researchers in swiftly and effectively extracting pertinent information from clinical notes, our work demonstrates the practical applicability of LLMs in the medical domain, particularly in information extraction from clinical notes and the development of chatbots utilizing this documentation. We highlight the significance of LLMs in assisting users to swiftly and effectively extract pertinent information from clinical notes.
The following are our contributions:
@@
  • We developed a conversational question answering system based on LLMs that allows natural language interaction for exploring clinical notes
@@
  • We conducted experiments evaluating combinations of different EM and LLMs to enhance performance and speed
@@
  • We demonstrated the advantages of weight quantization and the superiority of retrieval augmented generation over domain-specific fine tuning
@@
Organization: The paper is organized as follows: We discuss related works in section 2. Section 3 describes the approach. Details of the dataset are available in section 4. Evaluation and results are presented in section 5. We discuss the limitations and future work in section 6 and section 7 concludes the paper.
2. Related Work
2.1. LLMs in information retrieval
@@
The intersection of LLMs and Information Retrieval (IR) has emerged as a pivotal area of research, transforming traditional approaches and reshaping the landscape of information access.
The foundational work of Zhu et al. (Zhu et al., 2023) offers a comprehensive survey covering diverse applications of LLMs in information retrieval (IR). This work spans various facets of IR, including query rewriting, retrieval, re-ranking, and reading. Additionally, an influential contribution in the information retrieval domain is the tutorial by Begoli et al. (Begoli et al., 2022) . This tutorial advances both theoretical and practical aspects of transformer-based approaches, emphasizing efficiency and robustness in large-scale information retrieval scenarios. Its goal is to provide the audience with insights gained from developing and fine-tuning transformer/BERT-based models for IR tasks.
@@
In contrast, our approach focuses on the utilization of off-the-shelf models, enabling the direct retrieval of information from provided resources without the need for intricate fine-tuning procedures. This streamlined method significantly boosts usability and efficiency, circumventing the time and computational expenses associated with fine-tuning. Furthermore, our research demonstrates that in our case the performance achieved through our direct utilization method surpasses that attained through domain-specific fine-tuning processes.
2.2. Medical question answering using LLMs
In recent years, the field of medical question answering has experienced remarkable advancements, fueled by the integration of LLMs. Various research works have significantly contributed to the pursuit of expert-level performance in this challenging domain. Lee et al. (Lee et al., 2019) addresses the unique challenges of biomedical text mining by introducing BioBERT, a domain-specific language representation model pre-trained on large-scale biomedical corpora. Outperforming generic models and prior state-of-the-art approaches, BioBERT excels in tasks such as biomedical named entity recognition, relation extraction, and question answering. Another impactful work by Chakraborty et al. (Chakraborty et al., 2020) emphasizes the urgency of advancing biomedical research amid the SARS-CoV-2 pandemic. BioMedBERT, a pre-trained model on the BREATHE v1.0 dataset, demonstrates exceptional efficacy in Question-Answering (QA) and Information Retrieval (IR) tasks. Achieving state-of-the-art results on BioASQ datasets, it exhibits superior performance when coupled with Elasticsearch for IR tasks. Additionally, Singhal et al. (Singhal et al., 2023) introduces Med-PaLM 2, a groundbreaking development in medical question answering. Addressing inherent challenges in prior models, Med-PaLM 2 leverages improvements in base LLMs, medical domain fine-tuning, and innovative prompting strategies, including a novel ensemble refinement approach. The model achieves an impressive score of up to 86.5% on the MedQA dataset, surpassing its predecessor by over 19% and establishing a new state-of-the-art in expert-level medical question answering.
@@ 2.2. Medical question answering using LLMs
Our work diverges by placing a specific emphasis on the integration of retrieval augmented generation (RAG) for medical question answering. While the aforemnetioned models, have made significant strides in achieving expert-level performance through advancements in base LLMs, domain fine-tuning, and innovative prompting strategies, our approach focusing on the integration of RAG for clinical notes aims to strike a balance between performance and resource efficiency, making the model more accessible for widespread adoption in medical settings.
2.3. Interplay Between Embedding Models and Language Models
@@
The exploration of embedding models (EM) and their impact on the output of LLMs represents an under-explored but crucial aspect in natural language processing (NLP). While many studies concentrate on the applications and performance of LLMs, the specific influence of EM on their output remains a relatively uncharted territory.
@@
In the work by Asudani et al. (Asudani et al., 2023), the authors delve into the application of word embedding models and deep learning models within the context of text analytics tasks. Conducting a systematic literature review, they classify and analyze relevant articles, emphasizing the growing interest in utilizing these techniques for analysis and prediction. The study categorizes literature based on text analytics applications, covering areas such as text classification, sentiment analysis, named entity recognition (NER), recommendation systems, biomedical text mining, and topic modeling. The paper concludes that the use of domain-specific word embeddings and the LSTM model can significantly enhance overall text analytics task performance. In our investigation, we shift our focus to a distinct domain that remains unexplored in the aforementioned study — medical question answering. Moreover, we explore a different suite of EM and language models, addressing areas that previous works did not specifically cover in their research.
The systematic investigations into emotion analysis using deep learning Xu et al. (Xu et al., 2020), text classification employing deep learning techniques Dogru et al. (Dogru et al., 2021), and a survey on the training and evaluation of word embeddings Torregrossa et al. (Torregrossa et al., 2021) revolve around evaluating the performance of word embedding and deep learning models within specific domains. These studies not only compare the efficacy of these models but also offer an overview of alternative approaches applied to analogous tasks. In our investigation, we explore the unexplored domain of medical question answering, employing a different suite of EMs and LLMs not specifically addressed in previous research.
2.4. Retrieval Augmented Generation (RAG) in Various Contexts
Retrieval Augmented Generation (RAG) has emerged as a transformative technique across diverse domains, revolutionizing information processing and decision-making processes.
In a work by Datta et al. (Datta et al., 2023), the authors propose two models, MAKG (Medical Appropriateness Knowledge Graph) and RAG-GPT (Retrieval Augmented Generation – Generative Pretrained Transformer). MAKG functions as an autonomous coarse-grained medical-inappropriateness vigilance model for payers and regulators, while RAG-GPT operates as a fine-grained LLM with human-in-the-loop for assessing medical appropriateness and inappropriateness.
Another work by Lewis et al. (Lewis et al., 2021) introduces RAG models that combine parametric and non-parametric memory components to enhance sequence-to-sequence (seq2seq) models. By endowing pre-trained, parametric-memory generation models with a non-parametric memory, RAG models achieve state-of-the-art results in open-domain extractive question answering and knowledge-intensive generation tasks.
In the same vein of the previous works, our research focuses on RAG for medical question answering and exploring how RAG in this context is more efficient than traditional models fine-tuning.
3. Approach
Figure 1. Information retrieval pipeline
3.1. Methodology
@@
Our approach primarily utilizes Retrieval augmented generation (RAG) (Lewis et al., 2021), a technique that merges the power of pre-trained language models with the benefits of external knowledge sources. RAG employs a two-step process: a retrieval step that fetches relevant documents from a vast corpus of text, and a generation step that crafts responses based on these retrieved documents and the given input query.
@@
This technique is particularly advantageous when the required response includes information that the LLM has not encountered during its training. By combining the ability to generalize from language model pretraining with the capacity to retrieve specific details from external sources, RAG can generate highly detailed and accurate responses while avoiding the overhead of fine tuning the language model on specific datasets.
@@
    • Content Splitting: The loaded content is then split into chuncks. This step is crucial as it breaks down the text into smaller units, to comply with the EM context size limitations. Throughout this step, we intentionally allow for some overlap between the different chunks. This strategy helps to maintain the contextual integrity of the data.
    • Text Embedding: Each chunk of text is then transformed into an embedding. These embeddings are numerical representations that encapsulate the semantic meaning of the text. The transformation process is carried out by an embedding model.
    • Vector Store: The embeddings, along with reference to the original content the embedding was created from, are subsequently stored in a vector database, which forms the backbone of our retrieval system.
  2. (2)
@@
    • Output Generation: The re-ranked vectors, which represent the selected documents, along with the input prompt, are passed as context to a LLM. This model is responsible for text summarization and final output generation.
Table 1. Embedding vs Language models
@@
We leveraged the following components of the LangChain framework (Chase, 2022) to achieve the previously described pipeline:
@@
Document Store: Efficiently stores and retrieves documents, enabling the creation of a robust retrieval system, a crucial aspect of the RAG model. It achieves this by using a vector database to store embeddings and references to the original content, providing a fast and scalable solution for handling large volumes of data.
Retriever: Utilizes an index to scan the vector store. By employing cosine similarity, it identifies and returns the most relevant documents to support the RAG model’s operations.
Memory: Stores information about past interactions, acting as a conversation history. This not only improves the quality of responses but also ensures a more natural and human-like conversational experience for users.
@@
Chain: Allows for the creation of coherent workflows by linking multiple components together. For example, a chain can be created to take user input, format it with a specific template, and then pass the formatted response to the language model for generation, enhancing the system’s versatility and functionality.
3.2. Constraints
Our approach encountered certain constraints that influenced the scope of our study. The challenges we faced include:
4. Dataset
The Medical Information Mart for Intensive Care MIMIC-IV (Johnson et al., 2023) , is a single-center dataset encompassing information about over 40k patients admitted to intensive care units (ICUs) at at the Beth Israel Deaconess Medical Center (BIDMC). This dataset includes various patient ICU information like medications and measurements. MIMIC-IV-Note, a subset of the MIMIC-IV dataset contains deindentified clinical notes for patients. Table 2 provides a high level statistics of the MIMIC-IV dataset.
Table 2. Dataset Statistics